5 The length (in centimeters) of the leaf of a certain plant is a continuous random variable with probability density functiocm. The standard deviation of the distribution is (Round to two decimal places as needed.)The probability that the random variable is between the mean and 1 standard deviation above the mean is (Round to two decimal
Answer #1

Solution The given pdf is free) = 5 14 for 3<x<5.8 and we have to find @ mean S.D. and P (42X341+) :: mean = f(x) = S FCN),koS. DIX) = 0.8083 * Now plusxalto) = P(4.4 <x<5.2083) 5.2.083 S.2083 5 Ta s [x dae = 4.4 14 4.4 11 S [ 5.2083 -6,4] pluLX Slut